# A Framework for Quantum Advantage

### v1

@@ Abstract
As quantum computing approaches the threshold where certain tasks demonstrably outpace their classical machines, the need for a precise, clear, consensus-driven definition of quantum advantage becomes essential. Rapid progress in the field has blurred this term across companies, architectures, and application domains. Here, we aim to articulate an operational definition for quantum advantage that is both platform-agnostic and empirically verifiable. Building on this framework, we highlight the algorithmic families most likely to achieve early advantage. Finally, we outline our vision for the near future, in which quantum computers enhance existing high-performance computing platforms, enabling new frontiers in chemistry, materials discovery, optimization, and beyond.
@@ I Introduction
Quantum computing has long held promise in solving computational problems that are intractable for classical computers [1, 2, 3, 4, 5, 6, 7]. With the astonishing improvement in hardware, software and quantum algorithms, this promise feels closer than ever before, e.g., as seen in several demonstrations that have shown how quantum processors are now capable of performing accurate computations at a scale that defies brute-force classical simulation of quantum circuits [8, 9]. As these capabilities continue to mature, a central question has come to the forefront: When will quantum computers deliver a tangible advantage in solving practical computational tasks?
This milestone, commonly referred to as quantum advantage, represents a pivotal threshold for the field. It serves not only as a benchmark for quantum technological progress, but also as a guiding framework for allocating research investments and identifying application domains with near-term impact potential. Establishing criteria for when and how quantum advantage is achieved is therefore essential for advancing both our scientific understanding and the deployment of quantum technologies in real-world settings.
In this work, we outline the general definition of quantum advantage and the criteria that are necessary to claim it, as well as the necessary elements to achieve it: methods of running accurate quantum circuits, HPC + quantum infrastructure, and performant quantum hardware. It is also crucial to focus our attention on algorithms that we know can be run reliably and have verifiably correct outputs, thus, we also outline the collection of algorithms that fit this description. Once quantum advantage has been attained in experimental settings, we envision that quantum computers will start enhancing existing HPC platforms, unlocking new frontiers in chemistry, materials science, optimization, and beyond.
@@ II Definition of Quantum Advantage
Quantum advantage denotes the execution of an information-processing task on quantum hardware that satisfies two essential criteria: i) the correctness of the output can be rigorously validated, and ii) it is performed with a demonstrably superior efficiency, cost-effectiveness, or accuracy than is attainable with classical computation alone. Although the concept appears straightforward, these criteria remain insufficiently articulated in the current literature- e.g. in order to have confidence in the correctness of the output produced by a quantum computer, one must demonstrably validate that the quantum computation is trustworthy. As we argue in the manuscript, this is primarily achieved through rigorous error bars as can be obtained from error correction or formally proven error mitigation methods. The result can also be validated if it is efficiently verifiable on a classical computer. Given the accelerated pace of progress in the field, a consensus definition of quantum advantage is needed to enable rigorous evaluations of forthcoming claims.
@@ III Problems poised for quantum Advantage
In this section, we discuss the family of problems that are strong candidates to reach quantum advantage. Strong claims of advantage need to rely, at least, on one of two things: verifiability of the result, or trust in the underlying quantum hardware, algorithms and methods executing the computation. In the context of current quantum processors, this could present a substantial challenge since, while certain problems allow for direct classical verification of quantum outputs, many relevant problems of interest do not [10]. In these cases, confidence must be established through the indirect certification of the quantum computation itself. A fault tolerant quantum computer (FTQC) is one way to make this certification, however, rigorous methods are emerging where the full overhead of FTQC is not needed. This is a necessary precursor for quantum advantage, and has been deemed in the literature as quantum utility [8]: the ability of quantum devices to perform accurate computations at scales beyond exact brute-force classical simulations. Provided that quantum utility is established, the focus shifts to identifying computational tasks and algorithms where quantum solutions provide evidence supporting the claim of advantage.
In what follows, we discuss three key families of computational problems that currently stand out as promising avenues for realizing quantum advantage: i) problems that are solved via sampling algorithms, ii) problems that are verifiable via the variational principle, and iii) problems whose solution reduces to calculating expectation values of observables. We examine each of these approaches in greater detail, with the aim of elucidating their respective potential for demonstrating quantum advantage.
@@
Sampling algorithms use the information contained in a set of samples from probability distributions defined by quantum circuits. According to our current understanding of computational complexity, these algorithms are well poised to solve problems that are difficult for classical computers, such as integer factorization using Shor’s algorithm [2]. Here, however, the apparent need for error correction and fault tolerance represents a major challenge that we do not anticipate to be solved in the short term.
Another task that falls into the sampling category is random circuit sampling (RCS) [11]. RCS could offer a potential candidate for quantum advantage, but verifying the correctness of the solutions to these problems is classically intractable at scale. A natural way to certify such computations is to ensure that the method of producing the samples is free (or nearly free) of errors. To date, the only way we know how to achieve this is through error correction, which is why several experimental demonstrations [12, 13, 14] that claim quantum advantage have, in our view, not yet achieved this result. Attempts to explore sampling with quantum computers in the analog mode using neutral atoms [15] or superconducting qubits in the XY model [16], are promising from a physics perspective, but raise similar concerns regarding verification and robustness. In short, these demonstrations fail to certify the accuracy of the computation, which negates a fair comparison between quantum and classical methods [17]. Overall, given the lack of efficient verification, we do not believe that random circuit sampling is a fully satisfactory route toward quantum advantage.
@@
However, recent theoretical work [18] has introduced a promising variant of random circuit sampling (RCS) based on peaked random circuits. Such circuits are engineered to produce a specific bitstring with non-negligible probability, while maintaining random-like behavior otherwise. This structure opens new avenues for demonstrating verifiable quantum advantage, as the “peakedness” of the distribution can be detected by analyzing sampled outputs. Although recent results show that peakedness may, in some scenarios, be simulated in quasi-polynomial time [19], this insight also enhances our understanding of when and how quantum circuits exhibit a classically tractable structure. Furthermore, the possibility that classical algorithms might directly identify peakedness, without full simulation, offers valuable feedback for designing more robust quantum advantage protocols. Overall, peaked random circuits provide fertile ground for exploring new complexity-theoretic frontiers towards verifiable quantum advantage, but will need more rigorous determination on hardness to be clear advantage candidates.
Variational principle type of problems
@@
A family of computational problems that are promising candidates for achieving quantum advantage is those that obey the variational principle. One such example is finding the ground-state energies in certain physical systems. Computing ground-state energies is an important application at the root of many computational challenges in condensed matter and high-energy physics, chemistry, and materials science. The ground-state problem has an important implication in the search for quantum advantage. First, if the algorithm used satisfies the variational principle, approximate solutions can be classically ranked by accuracy, according to the energy approximations they produce. As a consequence, if a quantum solution has a better accuracy or lower energy, that would indicate a possible verifiable quantum advantage.
@@
Verifiability can occur at different levels depending on the nature of the algorithm. One can have variational quantum solutions that come with an uncertainty error, without it, or - for some algorithms - the entire solution proposed by the quantum computer can be classically representable. In the following, we discuss a few examples of quantum algorithms that have a range of verifiability levels.
@@
The quantum phase estimation algorithm (QPE) [20] is one of the first algorithms proposed for the ground-state problem. Provided with a good reference state with a polynomial overlap with the true ground state, QPE can solve the ground state problem efficiently in polynomial time. The solution from QPE algorithm comes with a precision error which is exponentially small in the size of its quantum auxiliary register, and its verifiability hinges on the trust that the quantum circuit is performed correctly and other approximations - such as those accompanying the Hamiltonian simulation part - are controlled. However, the quantum circuits that are necessary for the phase estimation algorithm are very deep, and they are generally thought to require fault-tolerant quantum computers.
@@
In the last decade, variational quantum eigensolvers (VQEs) [21, 22] and their adaptive versions [23] have been proposed as a low-depth alternative for ground-state problems, which can be executed on pre-fault-tolerant quantum devices. They can be seen as quantum equivalent of variational quantum Monte Carlo methods [24]. Compared to phase estimation, they do not have convergence guarantees, and are not directly tied to the quality of the initial reference state. Although the approximate estimations of the ground-state energies produced by VQE are variational, they come with an uncertainty related to the statistical nature of the energy estimation. This uncertainty allows the approximate solution to underestimate the ground state energy, breaking a strict ranking of the solution accuracy. However, one can take into account these uncertainties and introduce quality metrics that incorporate them, such as the V-score defined in [25]. While still constituting a good option for lattice models, low-depth VQE is not a good choice for utility-scale chemistry systems, since the number of measurements necessary to estimate molecular energy to useful precisions is prohibitive [26, 27].
The Krylov quantum diagonalization (KQD) [28], a quantum equivalent of classical algorithms based on power methods, bridges QPE and VQE, since it inherits convergence properties of QPE, and the variational behavior and statistical uncertainties of VQE. It removes the need to design a variational ansatz, and hinges on quantum circuits that simulate real-time evolutions. The verifiability of its solutions can be treated on the same level of VQE, and the solution accuracy can be ranked accordingly. Still, KQD cannot deal with large chemistry problems as a result of the same measurement problem that affects VQE.
@@
Recently, new algorithms designed to be executed on quantum-centric supercomputing platforms [9, 29] have overcome different issues that were affecting pre-fault-tolerant ground-state methods. Instead of estimating expectation values, as in VQE or KQD, the sample-based quantum diagonalization (SQD) uses measurements from a quantum ansatz state to identify a subspace onto which the original Hamiltonian can be projected. The projected Hamiltonian has a much lower dimension than the original and can be diagonalized using distributed classical computing. Remarkably, SQD has already been enabled to use quantum data to estimate energies for systems such as the iron sulfide cluster 2Fe-2S, matching classical state-of-the-art results, and tackled larger 4Fe-4S clusters beyond exact diagonalization [9]. The fact that the energies are estimated on a classical computer at machine precision allows one to estimate molecular energies that exceed what is achievable with VQE, at the same time dealing with errors from a quantum computer at classical machine precision, effectively removing the impact of useless quantum samples at the cost of a classical overhead.
@@
All the sample-based methods for ground states, both SQD and SKQD, achieve the highest degree of verifiability. The fact that the approximate ground states produced by these methods can be stored in classical memory allows for a high-precision ranking of their accuracy, without the need to factor in uncertainties coming from statistical noise. Quantum advantage obtained with these methods would therefore be classically certifiable and reproducible.
Quantum computers in the analog mode also allow one to generate bitstring distributions for regimes that can be challenging for classical methods. This includes regimes close to criticality or frustrated systems. Such examples are potential candidates for quantum-centric approaches such as SQD or SKQD.
Similarly to sample-based methods for quantum systems, the quantum approximate optimization algorithm (QAOA) outputs candidate solutions for combinatorial problems [32] that are classically verifiable. While we do not expect a general quantum speed-up for most optimization problems, we see potential for finding better solutions than classical solvers for specific problem instances with QAOA and other quantum heuristic approaches. To track progress towards potential quantum advantage and drive a fair and systematic benchmarking effort, a quantum optimization benchmark library has been introduced, with a curated set of 10 combinatorial optimization problem classes, corresponding problem instances, and well-defined reporting metrics [33]. Along similar lines, Cazals et al. [34] have identified hard instances that can be natively suited for neutral-atom-based analog quantum computers. Furthermore, neutral atoms arranged in amorphous arrays have been proven to have a spin-glass phase [35], which is expected to be hard to sample from by classical means.
Crucially, for many problem classes, instances already become challenging for state-of-the-art classical solvers at moderate problem sizes for known sparse model formulations. This indicates a strong compatibility with near-term quantum hardware. Nevertheless, based on the comparative strength of classical optimization methods, we believe that the most promising path forward is to target individual problem classes and even specific problem instances.
@@
The calculation of expectation values of observables of short-depth quantum circuits is an essential task for several near-term quantum algorithms targeting a range of computational problems: from estimating ground-state energies of molecules [21], measuring kernels in machine learning [36], to simulating the time evolution of spin correlation functions [8]. In the absence of fault tolerance, these estimations have also been shown to be biased by noise in early implementations of variational algorithms [22] with modest circuit depths. However, the invention of error mitigation methods, detailed later in Section IV, has provided a path to suppressing [37, 38] and even canceling [37] the effect of noise on obeservable estimation, albeit at exponential sampling cost. However, improvements in hardware error rates and sampling rates have scaled these methods to qubit and gate counts that are beyond brute-force classical computation [8] and could even outperform state-of-the-art classical approximations.
Simulating time evolution is a classic example of measuring expectation values of short-depth quantum circuits and a strong candidate for quantum advantage, that is important to consider from the lens of verifiability. Models with conserved quantities offer some verification of the quantum computation, but in the most general case, verifiability can be a challenge at scales beyond exact diagonalization. At such scales, as we will elaborate more in Section IV, it is particularly crucial to have error mitigation methods that can provide expectation values with rigorous error bars, to place trust in the result of the quantum computation.
Simulating time evolution is also particularly well suited for analog quantum simulators, and has already been studied with these systems at scales inaccessible to exact diagonalization, particularly for specific classes of Hamiltonians [39]. While controlling and characterizing errors in these systems remains an open challenge, recent theoretical results suggest that certain physically meaningful observables (e.g., local) can exhibit robustness to both deterministic and stochastic perturbations, with error bounds that may not scale with system size [40, 41, 42]. These insights, together with ongoing developments in analog error mitigation [43, 44], point to promising paths to enhance the reliability of analog quantum simulation.
Based on these algorithmic groups, several application domains are widely considered the most promising areas that could quickly take advantage of quantum advantage. Quantum simulations of chemistry and materials science [9, 45, 46, 47, 48] are widely regarded as one of the most promising near-term applications. These problems are often intractable for classical computers due to the exponential complexity of simulating quantum systems with high accuracy. Quantum computers, in contrast, can efficiently represent other quantum systems, and quantum algorithms exist to simulate both static and dynamical properties of quantum systems that have no efficient classical counterpart and can, in principle, lead to quantum advantages.

### v2

@@ Abstract
As quantum computers approach the threshold where they can demonstrably outpace their classical counterparts on certain tasks, the need for a precise, consensus-driven definition of quantum advantage becomes essential. Rapid progress in the field has blurred this term across organizations, architectures, and application domains. Here, we aim to articulate a functional definition for quantum advantage that is both platform-agnostic and empirically verifiable. Building on this framework, we highlight the algorithmic families most likely to achieve early advantage. Finally, we outline our vision for the near future, in which quantum computers enhance existing high-performance computing platforms.
@@ I Introduction
Quantum computing has long promised to solve computational problems that are intractable for classical computers [1, 2, 3, 4, 5, 6, 7]. Recent advances in quantum hardware, software, and algorithm design have significantly accelerated progress toward this goal. Notably, several experimental demonstrations have shown that quantum processors can now perform accurate computations at scales that surpass the capabilities of brute-force classical simulation of quantum circuits [8, 9]. As these technologies continue to mature, a critical question has emerged at the forefront of the field: When will quantum computers deliver a clear, demonstrable advantage in solving practical, real-world computational tasks? This milestone, commonly referred to as quantum advantage, represents a pivotal threshold for the field.
In this work, we propose a functional definition of quantum advantage and delineate the criteria required to substantiate such a claim. Establishing such criteria for when and how quantum advantage is achieved is crucial, not only as a rigorous benchmark of quantum technological progress, but also for identifying application domains with the potential for near-term impact. We also identify important components in this path: methods for the reliable execution of quantum circuits on performant quantum processors, their integration with high-performance computing (HPC) infrastructure, and the development of algorithms that can be executed with high fidelity and yield verifiably correct results.
@@ II Definition of Quantum Advantage
Quantum advantage denotes the execution of an information processing task on quantum hardware that satisfies two essential criteria: i) the correctness of the output can be rigorously validated, and ii) it is performed with a quantum separation that demonstrably offers superior efficiency, cost-effectiveness, or accuracy than what is attainable with classical computation alone. Although the concept appears straightforward, these criteria remain insufficiently articulated in the current literature.
II.1 Validation
Establishing trust in the outcome of a computation is one of the foremost tasks in establishing any new computational method. This is particularly critical for a quantum computation, since quantum computers are poised to solve problems that are not efficiently solvable by established classical methods, making a direct comparison challenging. Furthermore, since quantum computing is such a new paradigm, the concern is that simple sanity checks based on reasonable intuition may not be sufficient. Testing on simple cases that can be efficiently evaluated exactly by classical means may not capture the challenges the computation faces when addressing harder instances of the problem. As we argue in the manuscript, there are different ways in which we can ensure that the quantum computation can be validated to produce the correct answer even at a scale where establishing trust in the computation becomes challenging. The most straightforward way is to ensure that the computation itself was performed correctly by providing rigorous error bars, as can be obtained from a fault-tolerant quantum computer, or formally proven error mitigation methods. However, there are also interesting computational problems where the possibility of validation is directly possible by the nature of the answer.
1.
Error bars and error bounds A universal way to ensure that the computation can be trusted is to ensure that every step of the computation was performed accurately. The way to achieve this formally is through proven error bars. There are different methods that can ensure such error bars, the cardinal approach being fault-tolerant quantum computing, cf. Section IV.11. Other methods, such as rigorously proven quantum error mitigation methods, cf. Section IV.12, and post-selected error detection, cf.Section IV.13, are also able to provide rigorous error bars. However, compared with a fault-tolerant quantum computer, error mitigation and error detection methods have a cost that is more favorable only at smaller scales, cf. Section III, which is then clearly outperformed by fault-tolerant quantum computation at larger scales.
2.
Problems with efficient classical verification It is believed that for certain problems the answers can only be found efficiently on a quantum computer. But some of these problems are structured in such a way that, once found, these answers can be verified efficiently on a classical computer. Examples include sampling problems, cf. Section III, where we are promised samples from a distribution with a specific structure, such as peakedness [10], or algebraic problems such as determining the prime factors of an integer [2]. For these examples, we don’t need to rigorously check every step of the quantum computation, as we can directly verify the output is correct using a classical computer.
3.
Variational problems Similarly, there are problems for which it may not be possible to efficiently verify the solution on a classical computer directly, but we can easily score the quality of a solution and compare it with competing classical approaches. This is, for example, the case in variational problems, cf. Section III, such as in the estimation of the ground state energy of a molecule. In variational approaches, we are always assured that the reported number is an upper bound to the true energy value. This can happen without a statistical error [9] or in presence of it [11]. Hence, solutions can be scored based on which one gives the lowest estimate. The same is true for classical optimization problems, where the solution with the smallest cost function value is the best answer. This can be established without ensuring that every step of the computation was performed accurately.
Although a formal validation of the computation is the foremost goal in establishing a quantum advantage, it is important to recognize the high relevance that experimental sanity checks play in building trust in the device. In particular, formally proven results always rely on a set of initial assumptions, such as accurate noise models for the quantum hardware or detailed knowledge of the underlying architecture, that must themselves be verified. These assumptions are usually easier to verify than the full quantum computation on its own. Additionally, heuristic methods [12, 13, 8, 14] play an important role in evaluating the capabilities of quantum computers. While these methods are not formally proven, they are valuable in two ways: first, they help lay the groundwork for a rigorous error analysis; and second, they provide experimental evidence for empirical quantum separations, paving the way for more concrete claims of quantum advantage.
II.2 Quantum separations
Strictly speaking, it is not known whether a substantial, unconditional algorithmic separation between quantum computing and classical computing exists. Although this is considered extremely unlikely, it remains possible that the set of problems efficiently solvable by quantum computers could still be within the reach of classical methods. Unconditional separations have been proven in specific settings [15], but these typically do not lead to the substantial, potentially exponential differences in performance that quantum computers are poised to deliver. All currently known separations depend either on the best known classical algorithms, as in the case of Shor’s factoring algorithm [2], or on assumptions in complexity, as for example in ref. [16], cf. Section III. In practical terms, any claimed separation must withstand the test of time and be verified against both existing and emerging classical computational methods.
Therefore, once claims of an experimentally demonstrated quantum advantage begin to emerge, they must be evaluated on a common basis according to the criteria articulated in this article. Furthermore, quantum advantage is unlikely to materialize as a single, definitive milestone; instead, it should be understood as a sequence of increasingly robust demonstrations. Any claim of quantum advantage should be framed as a falsifiable scientific hypothesis, one that invites scrutiny, reproduction, and challenge from the broader research community. A cautious, iterative, and interpretive stance is therefore essential until state-of-the-art classical methods have had adequate opportunity to challenge the claim. However, should a classical approach later supersede a previously reported quantum advantage benchmark, the outcome should not be viewed as a failure; rather, it exemplifies constructive behavior that drives progress in both classical and quantum computation.
@@ III Problems poised for quantum Advantage
In what follows, we discuss three key families of computational problems that currently stand out as promising avenues for realizing quantum advantage: i) problems that are solved via sampling algorithms, ii) problems that are verifiable via the variational principle, and iii) problems whose solution reduces to calculating expectation values of observables. We examine each of these approaches in greater detail, with the aim of elucidating their respective potential for realizing quantum advantage.
@@
Sampling algorithms use all the information contained in a set of samples drawn from probability distributions produced by quantum circuits. Based on our current understanding of computational complexity theory, such algorithms are well positioned to address problems that are classically intractable — for example, integer factorization using Shor’s algorithm [2]. In this specific case, it is straightforward to verify that a proposed solution is correct through an efficient classical computation. However, the practical realization of these algorithms is hindered by the need for fault-tolerant quantum computation, which remains significant challenges that are unlikely to be resolved in the near term.
Similarly, the quantum approximate optimization algorithm (QAOA) outputs candidate solutions for combinatorial problems [17] that are classically verifiable. While we do not expect a general quantum speed-up for most optimization problems, we see potential for finding better solutions than classical solvers for specific problem instances with QAOA and other quantum heuristic approaches. To track progress towards potential quantum advantage and drive a fair and systematic benchmarking effort, a quantum optimization benchmark library has been introduced, with a curated set of 10 combinatorial optimization problem classes, corresponding problem instances, and well-defined reporting metrics [18]. Crucially, for many problem classes, instances have already become challenging for state-of-the-art classical solvers at moderate problem sizes for known sparse model formulations.
Along similar lines, Cazals et al. [19] have identified hard instances that can be natively suited for neutral-atom-based analog quantum computers. Furthermore, neutral atoms arranged in amorphous arrays have been proven to have a spin-glass phase [20], which is expected to be hard to sample from by classical means. This indicates a strong compatibility with near-term quantum processors. Nevertheless, based on the comparative strength of classical optimization methods, we believe that the most promising path forward is to target individual problem classes and even specific problem instances.
Another prominent example within the sampling paradigm is random circuit sampling (RCS) [16], which has been proposed as a potential candidate to demonstrate quantum advantage. Yet, verifying the correctness of RCS outputs at scale is classically intractable. A natural approach to certifying such computations involves ensuring that the sampling process is effectively error-free. To date, the only universally accepted method for achieving this is fault-tolerant quantum computing. Consequently, we argue that several experimental claims of quantum advantage [21, 22, 23] remain unsubstantiated under this criterion. Alternative strategies for ensuring that individual samples are drawn faithfully include error detection techniques based on post-selection [24] and single-shot error mitigation using coherent Pauli checks [25], cf. Section IV.13. Exploratory efforts in analog quantum sampling — such as those employing neutral atoms [26] or superconducting qubits in the XY model [27] — are promising from a physics perspective, but they similarly face challenges related to verification and robustness.
In short, these demonstrations have not succeeded in certifying the accuracy of the underlying computations, thereby precluding a fair comparison between quantum and classical methods [28]. Given the absence of efficient and scalable verification techniques, we conclude that random circuit sampling (RCS) does not yet constitute a fully satisfactory pathway to quantum advantage.
@@
However, recent theoretical developments [10] have introduced a promising variant of RCS based on peaked random circuits. These circuits are designed to generate a specific bitstring with non-negligible probability while preserving random-like characteristics across the remainder of the output distribution. This structure enables new opportunities for demonstrating verifiable quantum advantage, as the “peakedness” can be detected through an analysis of the sampled outputs. Although recent findings suggest that such peakedness may, in some cases, be simulated in quasi-polynomial time [29], this insight contributes to a deeper understanding of when and how quantum circuits exhibit classically tractable structure. Moreover, the possibility that classical algorithms might detect peakedness without requiring full simulation offers valuable guidance for designing more robust quantum advantage protocols. Overall, peaked random circuits provide fertile ground for exploring new complexity-theoretic frontiers towards verifiable quantum advantage, but will need more rigorous determination on hardness to be clear advantage candidates.
Variational problems
@@
A promising family of computational problems for achieving quantum advantage consists of those governed by the variational principle. A key example is the estimation of ground-state energies in quantum systems—a foundational task underlying many challenges in condensed matter physics, high-energy physics, quantum chemistry, and materials science. The ground-state problem is particularly relevant in the context of quantum advantage for several reasons. First, when the algorithm employed satisfies the variational principle, approximate solutions can be classically ranked by their energy estimates, even in the absence of an exact solution. Consequently, if a quantum solution yields a solution with lower energy than any classically obtained counterpart, it offers a potentially verifiable instance of quantum advantage.
@@
The degree of verifiability in quantum algorithms can vary significantly depending on the nature of the algorithm. One can have variational quantum solutions that come with/without an uncertainty error, or - for some algorithms - the entire solution proposed by the quantum computer can be classically representable. In the following sections, we examine several representative quantum algorithms, highlighting the spectrum of verifiability they offer and the implications for demonstrating quantum advantage.
@@
The Quantum Phase Estimation (QPE) algorithm [30] is one of the earliest quantum algorithms proposed for solving the ground-state energy problem. Given a suitable reference state with polynomial overlap with the true ground state, QPE can efficiently estimate the ground-state energy in polynomial time. The algorithm provides solutions with a precision error that decreases exponentially with the size of the quantum auxiliary register and its verifiability hinges on the trust that the quantum circuit is performed correctly and other approximations - such as those accompanying the Hamiltonian simulation part - are controlled. However, the quantum circuits that are necessary for the phase estimation algorithm are very deep and are generally thought to require fault-tolerant quantum computers.
@@
In the last decade, variational quantum eigensolvers (VQEs) [31, 32] and their adaptive versions [33] have been proposed as a low-depth alternative for ground-state problems, which can be executed on pre-fault-tolerant quantum devices. They can be seen as the quantum equivalent of variational quantum Monte Carlo methods [34]. Compared to phase estimation, they do not have convergence guarantees, and are not directly tied to the quality of the initial reference state. Although the approximate estimations of the ground-state energies produced by VQE are variational, they come with an uncertainty related to the statistical nature of the energy estimation. This uncertainty allows for the possibility of approximate solutions to underestimate the true ground-state energy, complicating a strict ranking of the solution accuracy when compared to other methods. However, quality metrics that incorporate uncertainty, such as the V-score introduced in [11]—have been proposed to better assess the quality of solutions. While still constituting a good option for lattice models, low-depth VQE is not a good choice for large-scale chemistry systems, since the number of measurements necessary to estimate molecular energy to useful precisions is prohibitive [35, 36].
Krylov Quantum Diagonalization (KQD) [37] is a recently proposed algorithm that bridges the gap between QPE and VQE. Inspired by classical power methods, KQD inherits the convergence properties of QPE while also exhibiting the variational behavior and statistical uncertainties characteristic of VQE. Unlike VQE, it does not require the design of a variational ansatz, relying instead on quantum circuits that simulate real-time evolution. The verifiability of KQD solutions is comparable to that of VQE, allowing the solution accuracy to be ranked on the basis of energy estimates and associated uncertainties. However, like VQE, KQD faces significant limitations when applied to large-scale quantum chemistry problems due to the prohibitive number of measurements required to achieve high-precision energy estimates.
@@
Recently, new quantum algorithms designed for execution on quantum-centric supercomputing platforms [9, 38] have addressed several limitations of pre-fault-tolerant ground-state methods. Instead of estimating expectation values, as in VQE or KQD, these algorithms uses measurements from a quantum ansatz state to identify a subspace onto which the original Hamiltonian can be projected. The projected Hamiltonian has a much lower dimension than the original Hamiltonian and can be diagonalized using distributed classical computing. Remarkably, Sample-based Quantum Diagonalization (SQD) has already been used to estimate energies for systems that are beyond exact diagonalization, such as the iron sulfide cluster 2Fe-2S with results that match classical state-of-the-art approximate methods such as density matrix renormalization group, and has been used to tackle larger 4Fe-4S clusters [9]. Since the final diagonalization is performed classically at machine precision, the resulting energy estimates are robust to quantum sampling errors, unlike VQE. Furthermore, the method mitigates the effect of errors in the quantum computer, removing the impact of useless quantum samples at the cost of classical overhead.
@@
Both the sample-based methods for ground states, SQD and SKQD, achieve the highest degree of verifiability. The fact that the approximate ground states produced by these methods can be stored in classical memory allows for a high-precision ranking of their accuracy, without the need to factor in uncertainties coming from statistical noise. Quantum advantage obtained with these methods would therefore be classically certifiable and reproducible. Quantum computers in the analog mode also allow one to generate bitstring distributions for regimes that can be challenging for classical methods. This includes regimes close to criticality or frustrated systems. Such examples are potential candidates for quantum-centric approaches such as SQD or SKQD.
@@
The calculation of expectation values of observables from short-depth quantum circuits is a fundamental task across a range of near-term quantum algorithms. These include applications such as estimating molecular ground state energies [31], evaluating quantum kernels in machine learning [41], and simulating time-dependent spin correlation functions [8]. In the absence of fault-tolerant hardware, early implementations of variational algorithms demonstrated that such estimations are susceptible to noise-induced bias, even at modest circuit depths [32]. The development of error mitigation techniques—discussed in detail in Section IV—has provided a pathway to suppress [42, 43] and, in some cases, cancel [42] the effects of noise on observable estimation. While these methods often incur an exponential sampling overhead, recent improvements in hardware fidelity and sampling rates have enabled their application to quantum circuits with qubit and gate counts beyond the reach of brute-force classical simulation [8]. In several cases, the resulting estimates are now competitive with state-of-the-art classical approximations.
Simulating time evolution is a classic example of measuring expectation values of short-depth quantum circuits and a strong candidate for quantum advantage, that is important to consider from the lens of verifiability. Models with conserved quantities offer some verification of the quantum computation, but in the most general case, verifiability can be a challenge at scales beyond exact diagonalization. At such scales, as we will elaborate more in Section IV, it is particularly crucial to have error mitigation methods that can provide expectation values with rigorous error bars, in order to place trust in the result of the quantum computation.
Simulating quantum time evolution is particularly well-suited to analog quantum simulators, which have already been used to explore many-body dynamics at scales beyond the reach of exact diagonalization for specific classes of Hamiltonians [44]. While the difficulty of achieving precise control and error characterization in these systems presents a general challenge for validating measured quantities, recent theoretical results suggest that certain physically meaningful observables—particularly local ones—can exhibit robustness to both deterministic and stochastic perturbations. Notably, the associated error bounds may remain independent of system size [45, 46, 47]. These insights, combined with ongoing progress in analog error mitigation techniques [48, 49], point to promising strategies for enhancing the reliability of analog quantum simulations.